# ReVision: High-Quality, Low-Cost Video Generation with Explicit 3D Physics Modeling for Complex Motion and Interaction

### v1

@@ Abstract
In recent years, video generation has seen significant advancements. However, challenges still persist in generating complex motions and interactions. To address these challenges, we introduce ReVision, a plug-and-play framework that explicitly integrates parameterized 3D physical knowledge into a pretrained conditional video generation model, significantly enhancing its ability to generate high-quality videos with complex motion and interactions. Specifically, ReVision consists of three stages. First, a video diffusion model is used to generate a coarse video. Next, we extract a set of 2D and 3D features from the coarse video to construct a 3D object-centric representation, which is then refined by our proposed parameterized physical prior model to produce an accurate 3D motion sequence. Finally, this refined motion sequence is fed back into the same video diffusion model as additional conditioning, enabling the generation of motion-consistent videos, even in scenarios involving complex actions and interactions. We validate the effectiveness of our approach on Stable Video Diffusion, where ReVision significantly improves motion fidelity and coherence. Remarkably, with only 1.5B parameters, it even outperforms a state-of-the-art video generation model with over 13B parameters on complex video generation by a substantial margin. Our results suggest that, by incorporating 3D physical knowledge, even a relatively small video diffusion model can generate complex motions and interactions with greater realism and controllability, offering a promising solution for physically plausible video generation.
Figure 1: By explicitly leveraging a parameterized 3D physical model, ReVision enhances pre-trained video generation models (e.g., Stable Video Diffusion [6]) to produce high-quality videos with complex motion (row 1), enabling precise motion control (rows 2, 3) and accurate interactions (rows 4, 5). During inference, an optional target pose can be specified through a rough sketch indicating the final position of a body part (blue circles in rows 1, 3, 4) or by a simple drag operation (blue arrows in row 2).
@@ 1 Introduction
Video diffusion models have recently achieved remarkable success in producing high-quality, temporally coherent videos [6, 7, 41, 24]. This progress has been driven by advances in model architectures [39], substantial increases in model complexity, now reaching tens of billions of parameters [41], and the availability of large-scale, high-quality video datasets [8]. However, current models still struggle to generate videos that adhere to realistic physical principles, making it difficult to consistently achieve fine-grained motion control, complex movements, and coherent object interactions. Despite extensive efforts to enhance performance through larger models and higher-quality datasets, a recent study [21] indicates that scaling model size and data alone is insufficient to fully capture the complexities of the physical world.
@@
On the other hand, human image animation models [16, 55, 67] offer valuable insights for addressing persistent challenges in video generation. Despite using smaller models and less data, these methods achieve consistent and precise video outputs with complex motions by following predefined 2D keypoint trajectories. This success suggests that incorporating a well-defined motion prior can substantially reduce the learning complexity of video generative models, enabling them to generate coherent and lifelike motion. However, in general video generation tasks, such strong guiding signals are typically unavailable, limiting the direct applicability of these animation techniques to broader video generation scenarios. This raises a critical question: Can we develop a video generation framework that leverages the implicit motion information embedded in the generated videos as guidance to further enhance video quality?
@@
In this paper, we propose a simple, plug-and-play video generation framework that explicitly optimizes physical knowledge in a conditional video generation model through a parameterized 3D representation, enabling the generation of complex object motions and interactions. The core of ReVision is to Regenerate Videos with explicit 3D physics representations, following an Extract–Optimize–Reinforce pipeline. Specifically, to effectively leverage 3D knowledge without heavy retraining of the diffusion model while preserving its original ability to generate high-quality visual appearance, we design the pipeline in three stages.
@@
In the first stage, we employ a video diffusion model (e.g., SVD [6]) to generate a coarse video conditioned on the given input. In the second stage, we utilize parametric 3D models (i.e., SMPL-X [37] for humans, SMAL [49, 80] for animals, and depth-aware 2D bounding boxes [69, 57] for objects) to extract 3D shape and motion features from the coarse video. These 3D representations are subsequently optimized by the proposed Parameterized Physical Prior Model (PPPM), producing a more accurate and natural 3D motion sequence. In the third stage, the refined 3D motion sequences are incorporated as additional conditioning inputs to reinforce the diffusion model, enabling it to regenerate the video with improved coherence and realism.
@@
Extensive qualitative results and human preference studies confirm that our model excels at generating complex motions and interactions. To demonstrate its applicability, we applied ReVision on Stable Video Diffusion (SVD) [6], substantially improving its ability to generate realistic and intricate motions. We further compare ReVision-SVD with HunyuanVideo [24], a state-of-the-art open-source video generation model with 13 billion parameters, and demonstrate superior motion quality. Finally, on the particularly challenging dance generation task, our model outperforms state-of-the-art human image animation methods that rely on ground-truth pose sequences, surpassing them across all evaluation metrics.
@@
  • We demonstrate that optimizing physical knowledge of the generative models enhances their ability to generate complex motions and interactions, suggesting a promising direction for improving video generation.
  • We introduce ReVision, a three-stage pipeline that significantly improves the motion quality of pre-trained video generation models by explicitly optimizing parameterized 3D physical knowledge extracted from generated videos.
  • We propose PPPM, a lightweight and robust parameterized physical prior model that effectively refines motion information in generated videos.
@@ 2 Related Work
Video Generation. With the success of diffusion models in image generation [46, 10, 27, 5], driven by advancements in both generative modeling strategies [13, 53, 26, 31, 28] and model architectures [3, 39, 29, 33], video generation [14, 52, 62, 70, 73, 77, 4, 41, 7] has recently attracted significant attention. Parallel to text-to-video (T2V) generation, image-to-video (I2V) methods [1, 25, 65, 36, 74] generate videos from a single starting frame. However, existing methods still struggle to handle complex motions and interactions, and often fail to maintain physical plausibility. To overcome these challenges, recent approaches incorporate additional conditions to enhance motion control in video generation. Common conditional inputs include text descriptions [17, 11, 9, 45, 72], which can further guide motion modeling. For example, MAGE [17] introduces a spatially aligned motion anchor to blend motion cues from text, and SEINE [9] uses a random-mask video diffusion model to create transitions guided by textual descriptions. Another popular condition is optical flow, where models [34, 35, 51] estimate rough flow from user-provided arrows or text to guide complex motion generation. In contrast, ReVision leverages implicit motion features already embedded in the generated video through 3D parameterized object representations. This allows it to directly extract, optimize, and reinforce accurate and reliable motion features from the generated video itself, resulting in precise motion sequences that enhance coherence and fidelity.
@@
Human Image Animation. Human image animation focuses on transferring motion from a source human to a target human by using ground-truth posture sequences, which can be represented as flow [63], keypoints [16, 55], or human part masks [67]. Extensive efforts have gone into extracting improved motion features. For example, MagicAnimate [67] leverages an off-the-shelf ControlNet [75] to obtain motion conditions, Hu et al. [16] introduce a Pose Guider network to align pose images with noise latents, and Animate-X [55] utilizes both implicit and explicit pose indicators to generate motion and pose features. Such strong guidance enables high-quality video generation in human image animation, as each posture sequence directly dictates the synthesis of corresponding frames. However, in general video generation, the ground-truth dense guidance is typically unavailable, and there is usually no reference video for extracting a motion sequence. To overcome this limitation, ReVision introduces a three-stage process: it first extracts an implicit, rough motion sequence from the generated video, then refines it using the proposed PPPM, and finally leverages the refined motion to guide video regeneration. This approach provides effective guidance for video generation, significantly improving video quality.
@@ 3 Preliminary
Figure 2: Method overview. Given the motion-conditioned video generation model, ReVision operates in three stages. Stage 1: A coarse video is generated based on the provided conditions (e.g., target pose, marked in blue, indicating the rough position of the yellow part in the last frame). Stage 2: 3D features from the generated coarse video are extracted and optimized using the proposed PPPM. Stage 3: The optimized 3D sequences are used to regenerate the video with enhanced motion consistency. Best viewed when zoomed in.
Latent Diffusion Model. Diffusion models [13] generate data through a denoising process that learns a probabilistic transformation. Latent diffusion models [46] move this process from pixel space to the latent space of a Variational Autoencoder (VAE) [22]. Specifically, we consider the latent representation $z_{0}$ of the input data. In the forward diffusion process, Gaussian noise is incrementally added to $z_{0}$:
@@
where $z_{t}$ represents the noisy latent representation at time step $t$, and $\gamma_{t}$ is a predefined noise schedule with $t\in(0,1)$. As $t$ increases, the cumulative noise applied to $z_{0}$ intensifies, gradually transforming $z_{t}$ closer to pure Gaussian noise. For simplicity, we can express the transformation from $z_{0}$ to $z_{t}$ directly as:
$z_{t}=\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\epsilon,$ (2)
where $\bar{\alpha}_{t}=\prod_{i=1}^{t}(1-\gamma_{i})$ and $\epsilon\sim\mathcal{N}(0,\mathbf{I})$. The latent diffusion model, parameterized by $\Theta$, learns to reverse this noising process by taking $z_{t}$ as input and reconstructing the clean data with the objective:
$\mathcal{L}=\left\|\epsilon-\epsilon_{\Theta}(z_{t},t,c)\right\|_{2}^{2},$ (3)
where $c$ is a condition (e.g., a text prompt) to guide the denoising process. Once the latent space is reconstructed, it is decoded via the VAE decoder.
@@
Video Latent Diffusion Model. We use SVD [6] as our base video diffusion model, which extends Stable Diffusion 2.1 [46] to video. The main architectural difference from image diffusion models is that SVD incorporates a temporal UNet [48] by adding temporal convolution and (cross-) attention [58] layers after each corresponding spatial layer.
@@
3D Human and Animal Mesh Recovery. We utilize the SMPL-X [37] and SMAL [79] parametric models to represent humans and animals, respectively. These models parameterize 3D meshes with pose parameters $\theta$ and shape parameters $\beta$. Additionally, SMPL-X model includes expression parameters $\psi$ to capture facial expressions through blend shapes. Given these parameters, SMPL-X model is a differentiable function that outputs a posed 3D human mesh $\mathcal{M}_{SMPL-X}(\theta_{h},\beta_{h},\psi_{h})\in\mathbb{R}^{10475\times 3}$, where pose $\theta_{h}\in\mathbb{R}^{165}$, shape $\beta_{h}\in\mathbb{R}^{10}$, and expression $\psi_{h}\in\mathbb{R}^{10}$. Similarly, SMAL model represents a posed 3D animal mesh with $\mathcal{M}_{SMAL}(\theta_{a},\beta_{a})\in\mathbb{R}^{3889\times 3}$, where pose $\theta_{a}\in\mathbb{R}^{105}$ and shape $\beta_{a}\in\mathbb{R}^{41}$. In this project, we recover 3D meshes by fitting SMPL-X and SMAL to our data and to the generated videos. This approach provides accurate part labeling and incorporates spatial priors of human and animal bodies, improving upon direct 2D shape predictions. Additionally, it enables motion strength computation by measuring movement speed in 3D space, offering higher accuracy than 2D pixel-based predictions.
@@
2.5D Parameterized Object Representation. Unlike humans and animals, there is no straightforward way to parameterize general objects in 3D space. Here, we represent objects in 2.5D by combining 2D bounding boxes [57], segmentation masks [40, 64], and depth estimation [69]. Specifically, given a detected bounding box and segmentation mask, we first convert the mask into a contour with 16 vertices. We then lift these 21 2D pixels (16 contour vertices, 4 bounding box corners, and the bounding box center) into 3D points using depth estimation. As a result, each object is now represented as a set of points $p_{o}\in\mathbb{R}^{21\times 3}$.
@@ 4 Method
ReVision requires extending a pre-trained video diffusion model to accept additional motion conditions as input. In Sec. 4.1, we describe how to adapt SVD into a motion-conditioned video generation model with minimal modifications. In Sec. 4.2, we introduce ReVision, a three-stage video generation pipeline built upon the extended SVD, incorporating a carefully designed Parameterized Physical Prior Model to provide accurate motion sequences as conditioning inputs. An illustration is provided in Fig. 2.
@@ 4.1 Motion-Conditioned Video Generation
Figure 3: Motion-conditioned video generation. We extend SVD for motion-conditioned video generation by introducing two additional conditioning channels to the original condition: (1) part segmentation mask derived from the 3D motion sequence, and (2) its corresponding confidence map.
@@
Note that all three settings use the same model architecture, introducing minimal modifications only to SVD’s first convolutional block. This design enables fine-tuning of only the initial convolutional block and all temporal layers, eliminating the need to train SVD from scratch. Consequently, this fine-tuning strategy allows the extended SVD to generate videos conditioned on various types of motion sequence inputs while preserving its ability to generate videos based only on text and the first frame.

### v2

# ReVision: Refining Video Diffusion with Explicit 3D Motion Modeling

@@ Abstract
In recent years, video generation has seen significant advancements. However, challenges still persist in generating complex motions and interactions. To address these challenges, we introduce ReVision, a plug-and-play framework that explicitly integrates parameterized 3D model knowledge into a pretrained conditional video generation model, significantly enhancing its ability to generate high-quality videos with complex motion and interactions. Specifically, ReVision consists of three stages. First, a video diffusion model is used to generate a coarse video. Next, we extract a set of 2D and 3D features from the coarse video to construct a 3D object-centric representation, which is then refined by our proposed parameterized motion prior model to produce an accurate 3D motion sequence. Finally, this refined motion sequence is fed back into the same video diffusion model as additional conditioning, enabling the generation of motion-consistent videos, even in scenarios involving complex actions and interactions. We validate the effectiveness of our approach on Stable Video Diffusion, where ReVision significantly improves motion fidelity and coherence. Remarkably, with only 1.5B parameters, it even outperforms a state-of-the-art video generation model with over 13B parameters on complex video generation by a substantial margin. Our results suggest that, by incorporating 3D motion knowledge, even a relatively small video diffusion model can generate complex motions and interactions with greater realism and controllability, offering a promising solution for physically plausible video generation. Project page: https://revision-video.github.io/
Figure 1: By explicitly leveraging a parameterized 3D motion model, ReVision enhances pre-trained video generation models (e.g., Stable Video Diffusion) to produce high-quality videos with complex motion (row 1), enabling precise motion control (rows 2, 3) and accurate interactions (rows 4, 5). During inference, an optional target pose can be specified via a rough sketch (rows 1, 3, 4, colored circles for different parts, dashed lines for the original pose, solid lines for the target pose) or a simple drag operation (blue arrows in row 2) indicating the final position.
@@ 1 Introduction
Video diffusion models have achieved remarkable success in producing high-quality, temporally coherent videos (Blattmann et al., 2023; Brooks et al., 2024; Polyak et al., 2024; Kong et al., 2024). It has been driven by advances in model architectures (Peebles & Xie, 2023), increases in model complexity, reaching tens of billion parameters (Polyak et al., 2024), and the availability of large-scale high-quality datasets (Chen et al., 2024). However, current models still struggle to generate videos that adhere to realistic physical principles, making it difficult to consistently achieve fine-grained motion control, complex movements, and coherent object interactions. Despite extensive efforts to improve performance through larger models and higher-quality datasets, a recent study (Kang et al., 2024) indicates that scaling model size and data alone is insufficient to fully capture the complexities of the real world.
@@
On the other hand, human image animation models (Hu, 2024; Tan et al., 2024; Xu et al., 2024) offer valuable insights for addressing persistent challenges in video generation. Despite using smaller models and less data, these methods achieve consistent and precise video outputs with complex motions by following predefined 2D keypoint trajectories. This success suggests that incorporating a well-defined motion prior can substantially reduce the learning complexity of video generative models, enabling them to generate coherent and lifelike motion. However, in general video generation tasks, such strong guiding signals are typically unavailable, limiting the direct applicability of these animation techniques to broader video generation scenarios. This raises a critical question: Can we develop a video generation framework that leverages the implicit motion information embedded in the generated videos as guidance to further enhance video quality?
@@
In this paper, we propose a simple, general, and plug-and-play video generation framework that incorporates motion knowledge into a conditional video generation model via a parameterized 3D representation, allowing the generation of videos with complex motions and interactions involving humans, animals, and general objects. The core of ReVision is to Regenerate Videos with explicit 3D motion representations, following an Extract–Optimize–Reinforce pipeline. Specifically, to effectively leverage 3D knowledge without heavy retraining of the diffusion model while preserving its original ability to generate high-quality visual appearance, we design the pipeline in three stages.
@@
In the first stage, we employ a video diffusion model, e.g., SVD (Blattmann et al., 2023), to generate a coarse video conditioned on the given input. In the second stage, we utilize parametric 3D models (i.e., SMPL-X (Pavlakos et al., 2019) for humans, SMAL (Rueegg et al., 2023; Zuffi et al., 2024) for animals, and 2D binary mask (Yu et al., 2023) with estimated depth (Yang et al., 2024b) for general objects) to extract 3D shape and motion features from the coarse video. These 3D object-centric representations are subsequently optimized by the proposed Parameterized Motion Prior model (PMP), producing a more accurate and natural 3D motion sequence. In the third stage, the refined 3D motion sequences are incorporated as additional conditioning inputs to reinforce the diffusion model, enabling it to regenerate the video with improved coherence and realism.
@@
Extensive qualitative results and human preference studies confirm that our model excels at generating complex motions and interactions. We first apply ReVision on Stable Video Diffusion (SVD) (Blattmann et al., 2023), substantially improving its ability to generate realistic and intricate motions. We further compare ReVision-SVD with HunyuanVideo (Kong et al., 2024), a state-of-the-art open-source video generation model with 13B parameters, and demonstrate superior motion quality. Finally, on the particularly challenging dance generation task, our model outperforms state-of-the-art human image animation methods that rely on ground-truth pose sequences, surpassing them across all evaluation metrics.
@@
  • We show that optimizing object-centric knowledge of generative models enhances their ability to generate complex motions and interactions, suggesting a promising direction for improving video generation.
  • We introduce ReVision, a three-stage pipeline that significantly improves the quality of pre-trained video generation models by explicitly optimizing parameterized 3D object-centric motion knowledge extracted from generated videos.
  • We propose PMP, a lightweight and robust parameterized motion prior model that effectively refines motion information in generated videos.
@@ 2 Related Work
Video Generation. With the success of diffusion models in image generation (Rombach et al., 2022; Esser et al., 2024; Liu et al., 2024a; Betker et al., 2023), driven by advancements in both generative modeling strategies (Ho et al., 2020; Song et al., 2020; Lipman et al., 2022; Liu et al., 2022; 2024b) and model architectures (Bao et al., 2023; Peebles & Xie, 2023; Liu et al., 2024c; Ma et al., 2024), video generation (Ho et al., 2022; Singer et al., 2022; Wang et al., 2024d; Yang et al., 2023; Zhang et al., 2024a; Zhou et al., 2022; Bar-Tal et al., 2024; Polyak et al., 2024; Brooks et al., 2024) has recently attracted significant attention. Parallel to text-to-video (T2V) generation, image-to-video (I2V) methods (Babaeizadeh et al., 2017; Li et al., 2018; Xiong et al., 2018; Pan et al., 2019; Zhang et al., 2020) generate videos from a single starting frame. However, existing methods still struggle to handle complex motions and interactions, and often fail to maintain physical plausibility. To overcome these challenges, recent approaches incorporate additional conditions to enhance motion control in video generation. Common conditional inputs include text descriptions (Hu et al., 2022; Girdhar et al., 2023; Chen et al., 2023; Ren et al., 2024b; Zeng et al., 2024), which can further guide motion modeling. For example, MAGE (Hu et al., 2022) introduces a spatially aligned motion anchor to blend motion cues from text, and SEINE (Chen et al., 2023) uses a random-mask video diffusion model to create transitions guided by textual descriptions. Another popular condition is optical flow, where models (Mahapatra & Kulkarni, 2022; Ni et al., 2023; Shi et al., 2024) estimate rough flow from user-provided arrows or text to guide complex motion generation. In contrast, ReVision leverages implicit motion features already embedded in the generated video through 3D parameterized object representations. This allows it to directly extract, optimize, and reinforce accurate and reliable motion features from the generated video itself, resulting in precise motion sequences that enhance coherence and fidelity.
@@
Human Image Animation. Human image animation focuses on transferring motion from a source human to a target human by using ground-truth posture sequences, which can be represented as flow (Wang et al., 2004), keypoints (Hu, 2024; Tan et al., 2024), or human part masks (Xu et al., 2024). Extensive efforts have gone into extracting improved motion features. For example, MagicAnimate (Xu et al., 2024) leverages an off-the-shelf ControlNet (Zhang et al., 2023a) to obtain motion conditions, Hu et al. (Hu, 2024) introduce a Pose Guider network to align pose images with noise latents, and Animate-X (Tan et al., 2024) utilizes both implicit and explicit pose indicators to generate motion and pose features. Such strong guidance enables high-quality video generation in human image animation, as each posture sequence directly dictates the synthesis of corresponding frames. However, in general video generation, the ground-truth dense guidance is typically unavailable, and there is usually no reference video for extracting a motion sequence. To overcome this limitation, ReVision introduces a three-stage process: it first extracts an implicit, rough motion sequence from the generated video, then refines it using the proposed PMP, and finally leverages the refined motion to guide video regeneration. This approach provides effective guidance for video generation, significantly improving video quality.
@@ 3 Preliminary
Latent Diffusion Model. Diffusion models (Ho et al., 2020) generate data through a denoising process that learns a probabilistic transformation. Latent diffusion models (Rombach et al., 2022) move this process from pixel space to the latent space of a Variational Autoencoder (Kingma & Welling, 2014). Specifically, we consider the latent representation $z_{0}$ of the input data. In the forward diffusion process, Gaussian noise is incrementally added to $z_{0}$:
@@
where $z_{t}$ represents the noisy latent representation at time step $t$, and $\gamma_{t}$ is a predefined noise schedule with $t\in(0,1)$. As $t$ increases, the cumulative noise applied to $z_{0}$ intensifies, gradually transforming $z_{t}$ closer to pure Gaussian noise. We express the transformation from $z_{0}$ to $z_{t}$ directly as: $z_{t}=\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\epsilon$, where $\bar{\alpha}_{t}=\prod_{i=1}^{t}(1-\gamma_{i})$ and $\epsilon\sim\mathcal{N}(0,\mathbf{I})$. The latent diffusion model, parameterized by $\Theta$, learns to reverse this noising process by taking $z_{t}$ as input and reconstructing the clean data with the objective: $\mathcal{L}=\left\|\epsilon-\epsilon_{\Theta}(z_{t},t,c)\right\|_{2}^{2}$, where $c$ is the condition to guide the denoising process. Once the latent space is reconstructed, it is decoded via the VAE decoder.
@@
Video Latent Diffusion Model. We use SVD (Blattmann et al., 2023) as our base video diffusion model, which extends Stable Diffusion 2.1 (Rombach et al., 2022) to video. The main architectural difference from image diffusion models is that SVD incorporates a temporal UNet (Ronneberger et al., 2015) by adding temporal convolution and (cross-) attention (Vaswani, 2017) layers after each corresponding spatial layer.
@@
3D Human and Animal Mesh Recovery. We utilize the SMPL-X (Pavlakos et al., 2019) and SMAL (Zuffi et al., 2017) parametric models to represent humans and animals, respectively. These models parameterize 3D meshes with pose parameters $\theta$ and shape parameters $\beta$. Additionally, SMPL-X model includes expression parameters $\psi$ to capture facial expressions through blend shapes. Given these parameters, SMPL-X model is a differentiable function that outputs a posed 3D human mesh $\mathcal{M}_{SMPL-X}(\theta_{h},\beta_{h},\psi_{h})\in\mathbb{R}^{10475\times 3}$, where pose $\theta_{h}\in\mathbb{R}^{165}$, shape $\beta_{h}\in\mathbb{R}^{10}$, and expression $\psi_{h}\in\mathbb{R}^{10}$. Similarly, SMAL model represents a posed 3D animal mesh with $\mathcal{M}_{SMAL}(\theta_{a},\beta_{a})\in\mathbb{R}^{3889\times 3}$, where pose $\theta_{a}\in\mathbb{R}^{105}$ and shape $\beta_{a}\in\mathbb{R}^{41}$. In our work, we recover 3D human and animal meshes by fitting the SMPL-X and SMAL models to both our data and the generated videos. This produces 3D mesh reconstructions for all humans and animals. Because these meshes are computer-graphics models with predefined body-part annotations at every vertex, we can obtain accurate part labels directly. The 3D meshes also allow us to compute motion strength by measuring movement speed in 3D space, which is more reliable than relying on 2D pixel motion alone.
@@
2.5D Parameterized Object Representation. Unlike humans and animals, there is no straightforward way to parameterize general objects in 3D space. Here, we represent objects in 2.5D by combining 2D bounding boxes (Varghese & Sambath, 2024), segmentation masks (Peng et al., 2020; Wu et al., 2024), and estimated depth (Yang et al., 2024b).
@@ 4 Method
Figure 2: Method overview. Given the video generation model, ReVision operates in three stages. Stage 1: A coarse video is generated based on the provided conditions (e.g., target pose, marked in blue, indicating the rough position of the yellow part in the last frame). Stage 2: 3D features from the generated coarse video are extracted and optimized using the proposed PMP. Stage 3: The optimized 3D sequences are used to regenerate the video with enhanced motion consistency. Best viewed when zoomed in.
ReVision requires extending a pre-trained video diffusion model to accept additional motion conditions as input. In Sec. 4.1, we describe how to adapt SVD into a motion-conditioned video generation model with minimal modifications. In Sec. 4.2, we introduce ReVision, a three-stage video generation pipeline built upon the extended SVD, incorporating a Parameterized Motion Prior model (PMP) to provide accurate motion sequences as conditioning inputs. An illustration is provided in Fig. 2.
@@ 4.1 Motion-Conditioned Video Generation
Figure 3: Motion-conditioned video generation. We enable motion-conditioned generation by introducing two extra conditioning channels: (1) part segmentation mask derived from the 3D motion sequence, and (2) its corresponding confidence map.
@@
Note that all three settings use the same model architecture, with minimal modifications limited to the first convolutional block of SVD. This design enables fine-tuning only the initial convolutional block and the temporal layers, avoiding the need to train SVD from scratch. As a result, the extended SVD can generate videos conditioned on various types of motion inputs, while still retaining its ability to generate videos from just text and the first frame.